# Nearsightedness in Materials with Indirect Band Gap

### v1

@@ 6. Proofs of the Main Results
$\nu_{0}(\bm{\xi}+i\bm{\zeta})\coloneqq\frac{\varepsilon_{N_{0}}(\bm{\xi})+\varepsilon_{N_{0}+1}(\bm{\xi})}{2},\qquad\bm{\xi},\bm{\zeta}\in\mathbb{R}^{d}$ (6.8)
@@
$\big{|}\nu(\bm{\xi}+i\bm{\zeta})-\nu_{0}(\bm{\xi}+i\bm{\zeta})\big{|}<\frac{\alpha}{2}\mathsf{gap}_{+},\quad\bm{\xi}+i\bm{\zeta}\in D_{\rm R}.$ (6.9)
@@
$\mathrm{dist}\left(\nu(\bm{\xi}+i\bm{\zeta}),\sigma\big{(}{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\big{)}\right)\geq\frac{\mathsf{g}(\xi)-3\alpha\mathsf{gap}_{+}}{2}\geq\frac{1-3\alpha}{2}\mathsf{g}(\bm{\xi}).$
@@
$C^{2}\coloneqq\int_{\Gamma^{\star}}|\widehat{\rho}(\bm{\xi}+i\bm{\zeta})_{\ell_{0}k_{0},ab}|^{2}\mathrm{d}\bm{\xi}=\sum_{k}\left|\rho^{\mathrm{ref}}_{\ell k,ab}\right|^{2}e^{2\,\bm{\zeta}\cdot(\ell-k)}<\infty.$ (6.10)
@@
$C^{2}\leq\int_{\Gamma^{\star}}\left(\frac{\ell(\mathscr{C})}{2\pi}\sup_{z\in\nu(\bm{\xi}+i\bm{\zeta})+\mathscr{C}}|R_{z}(\bm{\xi}+i\bm{\zeta})_{\ell_{0}k_{0},ab}|\right)^{2}\mathrm{d}\bm{\xi}\leq\left(\frac{2\ell(\mathscr{C})}{\pi}\right)^{2}\int_{\Gamma^{*}}\mathsf{g}(\bm{\xi})^{-2}{\,{\rm d}}\bm{\xi}$ (6.11)
where $\ell(\mathscr{C})$ denotes the length of the contour. Here, we have used the fact that $|R_{z}(\bm{\xi})_{\ell_{0}k_{0},ab}|\leq\mathrm{dist}\big{(}z,\sigma({\mathcal{H}}_{\boldsymbol{\xi}})\big{)}^{-1}\leq 4\mathsf{g}(\bm{\xi})^{-1}$. ∎
6.2. Proof of Section 3.1
@@
$\Big{|}\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}_{\ell k,ab}\Big{|}\leq\frac{4}{\mathfrak{d}^{\mathrm{ref}}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})r_{\ell k}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})(|\ell|+|k|)},$ (6.12)
@@
$C(u)$ $=CR^{d}_{1}e^{\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{1}}\frac{\|Du\|_{\ell^{2}_{\Upsilon}}}{(\mathfrak{d}^{\rm ref})^{d+4}\big{(}\mathfrak{d}(u)\big{)}^{d+1}},$ (6.13)
where $c_{0},C>0$ depend only on $d,\mathfrak{m},h_{0},\gamma_{0},N_{\mathrm{b}},M$, and $R_{1}>0$ is sufficiently large such that $\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\setminus B_{R_{1}})}\leq c_{1}\mathfrak{d}^{\mathrm{ref}}$ for some $c_{1}>0$. In particular, for sufficiently small $\|Du\|_{\ell^{2}_{\Upsilon}}$, $R_{1}$ can be chosen arbitrarily.
@@
The point is mainly based on the approximation of ${\mathcal{H}}(u)$ by the finite rank update of ${\mathcal{H}}^{\mathrm{ref}}$ [27]. We summarise the proof here, as parts of the ideas will be revisited in the subsequent discussion. [27, Lemma 5.4] states that, for each $\delta>0$, there exist $R_{\delta}>0$ (depending on $\delta$) and operators $P_{\delta}(u),P_{\rm{loc}}(u)$ such that
${\mathcal{H}}(u)={\mathcal{H}}^{\mathrm{ref}}+P_{\delta}(u)+P_{\rm{loc}}(u),$ (6.14)
where $\|P_{\delta}(u)\|_{\mathrm{F}}\leq\delta$, and $[P_{\rm{loc}}(u)]^{ab}_{\ell k}=[{\mathcal{H}}(u)-{\mathcal{H}}^{\rm ref}]_{\ell k}^{ab}$ if $(\ell,k)\in B_{R_{\delta}}\times B_{R_{\delta}}$; else, $[P_{\rm{loc}}(u)]^{ab}_{\ell k}=0$. That is, we can decompose ${\mathcal{H}}(u)-{\mathcal{H}}^{\rm ref}$ into two parts: $P_{\delta}$ and $P_{\rm loc}$, which are small in terms of Frobenius norm and rank, respectively. As for the selection of $R_{\delta}$, by [27, Lemma 5.4], we can first choose a sufficiently large $R_{1}>0$ satisfying $|u(\ell)-u(k)|\leq\mathfrak{m}|\ell-k|$ for all $\ell,k\in\Lambda\setminus B_{R_{1}}$, so that
$\Big{|}{\mathcal{H}}(u)_{\ell k,ab}-{\mathcal{H}}^{\mathrm{ref}}_{\ell k,ab}\Big{|}$ $\leq h_{0}e^{-\frac{\sqrt{3}}{2}\gamma_{0}\mathfrak{m}r_{\ell k}}|u(\ell)-u(k)|\quad\forall\ell,k\in\Lambda\setminus B_{R_{1}},$ (6.15)
and then, for $\forall R_{2}>0$, the regularity of ${\mathcal{H}}(u)$ and ${\mathcal{H}}^{\mathrm{ref}}$ gives
$\sum_{\ell\in B_{R_{1}}}\sum_{\begin{subarray}{c}k\in\Lambda\backslash B_{R_{1}}\\ {|\ell-k|>R_{2}}\end{subarray}}\Big{|}{\mathcal{H}}(u)_{\ell k,ab}-{\mathcal{H}}^{\mathrm{ref}}_{\ell k,ab}\Big{|}^{2}\leq C\sum_{\ell\in B_{R_{1}}}\sum_{\begin{subarray}{c}k\in\Lambda\backslash B_{R_{1}}\\ {|\ell-k|>R_{2}}\end{subarray}}e^{-2\gamma_{0}\mathfrak{m}|\ell-k|}\leq CR_{1}^{d}e^{-\gamma_{0}\mathfrak{m}R_{2}},$ (6.16)
[…]
@@
$\leq C\big{(}\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R_{1}})}^{2}+R_{1}^{d}e^{-\gamma_{0}\mathfrak{m}R_{2}}\big{)},$ (6.17)
@@
$C\big{(}\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R_{1}})}^{2}+R_{1}^{d}e^{-\gamma_{0}\mathfrak{m}R_{2}}\big{)}\leq\big{(}\frac{1}{2}\mathfrak{d}^{\mathrm{ref}}\big{)}^{2},$ (6.18)
@@
$\begin{split}&\quad\left|\left[U\big{(}I-VA^{-1}U\big{)}^{-1}V\right]_{\ell k,ab}\right|=\left|\left[A\big{(}A-P_{\rm loc}(u)\big{)}^{-1}P_{\rm loc}(u)\right]_{\ell k,ab}\right|\\ &\leq\sum_{\begin{subarray}{c}m,n:|\bm{r}_{n}|\leq R_{\delta}\end{subarray}}\sum_{a_{1}b_{1}}\Big{|}\big{(}z-{\mathcal{H}}^{\mathrm{ref}}-P_{\delta}(u)\big{)}_{\ell m,aa_{1}}\Big{|}\left|\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}_{mn,a_{1}b_{1}}\right|\big{|}P_{\rm loc}(u)_{nk,b_{1}b}\big{|}\\ &\leq\frac{2N_{\mathrm{b}}(|z|+h_{0})}{\mathfrak{d}(u)}\left(\sum_{\begin{subarray}{c}m,n:|\bm{r}_{n}|\leq R_{\delta}\end{subarray}}e^{-\gamma_{0}r_{\ell m}}e^{-\gamma_{\rm{CT}}(\mathfrak{d}(u))r_{mn}}\right)\max_{\begin{subarray}{c}\ell,k\in B_{R_{\delta}}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\big{|}P_{\mathrm{loc}}(u)_{\ell k,ab}\big{|}\\ &\leq\frac{c}{\big{(}\mathfrak{d}(u)\big{)}^{d+1}}\max_{\begin{subarray}{c}\ell,k\in B_{R_{\delta}}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\big{|}P_{\mathrm{loc}}(u)_{\ell k,ab}\big{|}\end{split}$ (6.20)
@@
$\leq\max_{\begin{subarray}{c}\ell,k\in B_{R_{\delta}}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\Big{|}\nabla h^{ab}_{\ell k}\big{(}\boldsymbol{\nu}_{\theta}\big{)}\cdot\big{[}u(\ell)-u(k)\big{]}\Big{|}\leq C\|Du\|_{\ell^{2}_{\Upsilon}(B_{R_{\delta}})},$ (6.21)
@@
$\leq C(u)\,e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})(|\ell|+|k|)},$ (6.22)
where $C(u):=Ce^{\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{\delta}}\|Du\|_{\ell^{2}_{\Upsilon}}\big{/}\big{(}(\mathfrak{d}^{\rm ref})^{d+2}\big{(}\mathfrak{d}(u)\big{)}^{d+1}\big{)}$ with $C$ depending on $h_{0},\gamma_{0},\mathfrak{m},d,N_{\mathrm{b}},M$. Noticing that, from (6.18), we can find some constants $C_{1},C_{2}>0$ such that $e^{\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{2}}=C_{1}R_{1}^{d}(\mathfrak{d}^{\mathrm{ref}})^{-2}$, and $\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\setminus B_{R_{1}})}<C_{2}\mathfrak{d}^{\mathrm{ref}}$. Hence, the prefactor $C(u)$ can be written as
$C(u)$ $=CR_{1}^{d}e^{\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{1}}\frac{\|Du\|_{\ell^{2}_{\Upsilon}}}{(\mathfrak{d}^{\rm ref})^{d+4}\big{(}\mathfrak{d}(u)\big{)}^{d+1}}.$
@@
$\begin{split}\left|\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}_{\ell k,ab}\right|&\leq\Big{|}A^{-1}_{\ell k,ab}\Big{|}+\left|\big{[}A^{-1}U\big{(}I-VA^{-1}U\big{)}^{-1}VA^{-1}\big{]}_{\ell k,ab}\right|\\ &\leq\frac{4}{\mathfrak{d}^{\mathrm{ref}}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})(|\ell|+|k|)}.\end{split}$ (6.23)
@@
$\left|\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}_{\ell k,ab}\right|\leq\left(\frac{4}{\mathfrak{d}^{\mathrm{ref}}}e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}+C(u)\right)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}:=\tilde{C}(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}.$ (6.24)
Additionally, we can always choose $\mathscr{C}$ in the above estimation satisfying
@@
With the above preparation, we now proceed to the proof of Section 3.1:
Proof of Section 3.1.
@@
$\big{|}\rho(u)_{\ell k,ab}\big{|}\leq\big{|}\rho^{\mathrm{ref}}_{\ell k,ab}\big{|}+\big{|}\rho(u)_{\ell k,ab}-\rho^{\mathrm{ref}}_{\ell k,ab}\big{|}\quad\forall\ell,k\in\Lambda,1\leq a,b\leq N_{\mathrm{b}}.$ (6.25)
To obtain the result, it suffices to only estimate the last term in (6.25), for the control of the first term has been given by Section 3.1. The difference between $\rho(u)$ and $\rho^{\mathrm{ref}}$ is bounded by

### v2

@@ 6.1. Proof of Section 3
(6.8) $\nu_{0}(\bm{\xi}+i\bm{\zeta})\coloneqq\frac{\varepsilon_{N_{0}}(\bm{\xi})+\varepsilon_{N_{0}+1}(\bm{\xi})}{2},\quad\bm{\xi},\bm{\zeta}\in\mathbb{R}^{d}$
@@
(6.9) $\big|\nu(\bm{\xi}+i\bm{\zeta})-\nu_{0}(\bm{\xi}+i\bm{\zeta})\big|<\frac{\alpha}{2}\mathsf{gap}_{+},\quad\bm{\xi}+i\bm{\zeta}\in D_{\rm R}.$
@@
(6.10) $\mathrm{dist}\left(\nu(\bm{\xi}+i\bm{\zeta}),\sigma\big({\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\big)\right)\geq\frac{\mathsf{g}(\bm{\xi})-3\alpha\mathsf{gap}_{+}}{2}\geq\frac{1-3\alpha}{2}\mathsf{g}(\bm{\xi}).$
@@
(6.11) $C^{2}\coloneqq\int_{\Gamma^{\star}}|\widehat{\rho}(\bm{\xi}+i\bm{\zeta})_{\ell_{0}k_{0},ab}|^{2}\mathrm{d}\bm{\xi}=\sum_{k}\left|\rho^{\mathrm{ref}}_{\ell k,ab}\right|^{2}e^{2\,\bm{\zeta}\cdot(\ell-k)}<\infty.$
@@
(6.12) $C^{2}\leq\int_{\Gamma^{\star}}\left(\frac{\ell(\mathscr{C})}{2\pi}\sup_{z\in\nu(\bm{\xi}+i\bm{\zeta})+\mathscr{C}}|R_{z}(\bm{\xi}+i\bm{\zeta})_{\ell_{0}k_{0},ab}|\right)^{2}\mathrm{d}\bm{\xi}\leq\left(\frac{2\ell(\mathscr{C})}{\pi}\right)^{2}\int_{\Gamma^{*}}\mathsf{g}(\bm{\xi})^{-2}{\,{\rm d}}\bm{\xi}$
where $\ell(\mathscr{C})$ denotes the length of the contour. Here, we have used the fact that $|R_{z}(\bm{\xi})_{\ell_{0}k_{0},ab}|\leq\mathrm{dist}\big(z,\sigma({\mathcal{H}}_{\boldsymbol{\xi}})\big)^{-1}\leq 4\mathsf{g}(\bm{\xi})^{-1}$. This concludes the estimate (3.2). ∎
6.2. Proof of Section 3
@@
(6.13) $\Big|\big(z-{\mathcal{H}}(u)\big)^{-1}_{\ell k,ab}\Big|\leq\frac{4}{\mathfrak{d}^{\mathrm{ref}}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})r_{\ell k}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})(|\ell|+|k|)},$
@@
(6.14) $C(u)\coloneqq C\frac{e^{\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{1}}}{(\mathfrak{d}^{\rm ref})^{d+4}}\frac{\|Du\|_{\ell^{2}_{\Upsilon}}}{\mathfrak{d}(u)^{d+1}},\quad R_{1}>0\text{ s.t. }\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\setminus B_{R_{1}})}\leq c_{1}\mathfrak{d}^{\rm ref}$
for some $c_{1}>0$, and $c_{0},C>0$ depend only on $d,\mathfrak{m},h_{0},\gamma_{0},N_{\mathrm{b}}$ and $M$. In particular, for sufficiently small $\|Du\|_{\ell^{2}_{\Upsilon}}$, $R_{1}$ can be chosen arbitrarily.
@@
The point is mainly based on the approximation of ${\mathcal{H}}(u)$ by the finite rank update of ${\mathcal{H}}^{\mathrm{ref}}$ [28]. We summarise the proof here, as parts of the ideas will be revisited in the subsequent discussion. [28, Lemma 5.4] states that, for each $\delta>0$, there exist $R_{\delta}>0$ (depending on $\delta$) and operators $P_{\delta}(u),P_{\rm{loc}}(u)$ such that
(6.15) ${\mathcal{H}}(u)={\mathcal{H}}^{\mathrm{ref}}+P_{\delta}(u)+P_{\rm{loc}}(u),$
where $\|P_{\delta}(u)\|_{\mathrm{F}}\leq\delta$, and $P_{\rm{loc}}(u)]^{ab}_{\ell k}=[{\mathcal{H}}(u)-{\mathcal{H}}^{\rm ref}]_{\ell k}^{ab}$ if $(\ell,k)\in B_{R_{\delta}}\times B_{R_{\delta}}$; else, $P_{\rm{loc}}(u)]^{ab}_{\ell k}=0$. That is, we can decompose ${\mathcal{H}}(u)-{\mathcal{H}}^{\rm ref}$ into two parts: $P_{\delta}$ and $P_{\rm loc}$, which are small in terms of Frobenius norm and rank, respectively. As for the selection of $R_{\delta}$, by [28, Lemma 5.4], we can first choose a sufficiently large $R_{1}>0$ satisfying $|u(\ell)-u(k)|\leq\mathfrak{m}|\ell-k|$ for all $\ell,k\in\Lambda\setminus B_{R_{1}}$, so that
(6.16) $\Big|{\mathcal{H}}(u)_{\ell k,ab}-{\mathcal{H}}^{\mathrm{ref}}_{\ell k,ab}\Big|$ $\leq h_{0}e^{-\frac{\sqrt{3}}{2}\gamma_{0}\mathfrak{m}r_{\ell k}}|u(\ell)-u(k)|\quad\forall\ell,k\in\Lambda\setminus B_{R_{1}},$
and then, for $R_{2}>0$, the regularity of ${\mathcal{H}}(u)$ and ${\mathcal{H}}^{\mathrm{ref}}$ gives
(6.17) $\sum_{\ell\in B_{R_{1}}}\sum_{\begin{subarray}{c}k\in\Lambda\backslash B_{R_{1}}\\ {|\ell-k|>R_{2}}\end{subarray}}\Big|{\mathcal{H}}(u)_{\ell k,ab}-{\mathcal{H}}^{\mathrm{ref}}_{\ell k,ab}\Big|^{2}\leq C\sum_{\ell\in B_{R_{1}}}\sum_{\begin{subarray}{c}k\in\Lambda\backslash B_{R_{1}}\\ {|\ell-k|>R_{2}}\end{subarray}}e^{-2\gamma_{0}\mathfrak{m}|\ell-k|}\leq CR_{1}^{d}e^{-\gamma_{0}\mathfrak{m}R_{2}},$
where $C>0$ depends on $d,\mathfrak{m},\gamma_{0},h_{0}$. Through (6.16) and (6.17), we have
@@
(6.18) $\leq C\big(\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R_{1}})}^{2}+R_{1}^{d}e^{-\gamma_{0}\mathfrak{m}R_{2}}\big),$
@@
(6.19) $C\big(\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R_{1}})}^{2}+R_{1}^{d}e^{-\gamma_{0}\mathfrak{m}R_{2}}\big)\leq\big(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}}\big)^{2},$
@@
(6.21) $\begin{split}&\quad\left|\left[U\big(I-VA^{-1}U\big)^{-1}V\right]_{\ell k,ab}\right|=\left|\left[A\big(A-P_{\rm loc}(u)\big)^{-1}P_{\rm loc}(u)\right]_{\ell k,ab}\right|\\ &\leq\sum_{\begin{subarray}{c}m,n:|\bm{r}_{n}|\leq R_{\delta}\end{subarray}}\sum_{a_{1}b_{1}}\Big|\big(z-{\mathcal{H}}^{\mathrm{ref}}-P_{\delta}(u)\big)_{\ell m,aa_{1}}\Big|\left|\big(z-{\mathcal{H}}(u)\big)^{-1}_{mn,a_{1}b_{1}}\right|\big|P_{\rm loc}(u)_{nk,b_{1}b}\big|\\ &\leq\frac{2N_{\mathrm{b}}(|z|+h_{0})}{\mathfrak{d}(u)}\left(\sum_{\begin{subarray}{c}m,n:|\bm{r}_{n}|\leq R_{\delta}\end{subarray}}e^{-\gamma_{0}r_{\ell m}}e^{-\gamma_{\rm{CT}}(\mathfrak{d}(u))r_{mn}}\right)\max_{\begin{subarray}{c}\ell,k\in B_{R_{\delta}}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\big|P_{\mathrm{loc}}(u)_{\ell k,ab}\big|\\ &\leq\frac{c}{\big(\mathfrak{d}(u)\big)^{d+1}}\max_{\begin{subarray}{c}\ell,k\in B_{R_{\delta}}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\big|P_{\mathrm{loc}}(u)_{\ell k,ab}\big|,\end{split}$
@@
(6.22) $\leq\max_{\begin{subarray}{c}\ell,k\in B_{R_{\delta}}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\Big|\nabla h^{ab}_{\ell k}\big(\boldsymbol{\nu}_{\theta}\big)\cdot\big[u(\ell)-u(k)\big]\Big|\leq C\|Du\|_{\ell^{2}_{\Upsilon}(B_{R_{\delta}})},$
@@
(6.23) $\leq C(u)\,e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})(|\ell|+|k|)},$
where $C(u):=Ce^{\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{\delta}}\|Du\|_{\ell^{2}_{\Upsilon}}\big/\big((\mathfrak{d}^{\rm ref})^{d+2}\big(\mathfrak{d}(u)\big)^{d+1}\big)$ with $C$ depending on $h_{0},\gamma_{0},\mathfrak{m},d,N_{\mathrm{b}}$ and $M$. Noticing that, from (6.19), we can find some constants $C_{1},C_{2}>0$ such that $e^{\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{2}}=C_{1}R_{1}^{d}(\mathfrak{d}^{\mathrm{ref}})^{-2}$, and $\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\setminus B_{R_{1}})}<C_{2}\mathfrak{d}^{\mathrm{ref}}$. Hence, the prefactor $C(u)$ can be written as
$C(u)=C\frac{e^{\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})R_{1}}}{(\mathfrak{d}^{\rm ref})^{d+4}}\frac{\|Du\|_{\ell^{2}_{\Upsilon}}}{\mathfrak{d}(u)^{d+1}}.$
@@
(6.24) $\begin{split}\left|\big(z-{\mathcal{H}}(u)\big)^{-1}_{\ell k,ab}\right|&\leq\Big|A^{-1}_{\ell k,ab}\Big|+\left|\big[A^{-1}U\big(I-VA^{-1}U\big)^{-1}VA^{-1}\big]_{\ell k,ab}\right|\\ &\leq\frac{4}{\mathfrak{d}^{\mathrm{ref}}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})(|\ell|+|k|)}.\end{split}$
@@
$\left|\big(z-{\mathcal{H}}(u)\big)^{-1}_{\ell k,ab}\right|\leq$ $\left(\frac{4}{\mathfrak{d}^{\mathrm{ref}}}e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})(|\ell|+|k|-r_{\ell k})}\right)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}}$
(6.25) $:=$ $C_{\ell k}(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\rm ref})r_{\ell k}},$
where $C_{\ell k}(u)$ satisfies $C_{\ell k}(u)\lesssim\frac{1}{\mathfrak{d}^{\mathrm{ref}}}$ as $|\ell|+|k|-r_{\ell k}\to\infty$. Additionally, we can always choose $\mathscr{C}$ in the above estimation satisfying
@@
With the above preparation, we now proceed to the proof of Section 3:
Proof of Section 3.
@@
(6.26) $\big|\rho(u)_{\ell k,ab}\big|\leq\big|\rho^{\mathrm{ref}}_{\ell k,ab}\big|+\big|\rho(u)_{\ell k,ab}-\rho^{\mathrm{ref}}_{\ell k,ab}\big|\quad\forall\ell,k\in\Lambda,1\leq a,b\leq N_{\mathrm{b}}.$
To obtain the result, it suffices to only estimate the last term in (6.26), for the control of the first term has been given by Section 3. Applying the finite rank update (6.15), the difference between $\rho(u)$ and $\rho^{\mathrm{ref}}$ is bounded by